Im not sure how to figure out the answer. How would I work this? Determine if the ordered pair (0,2) is a solution of -3x + y = -2
Thanks a bunch
Kim

Answer by venugopalramana(3286)

Determine if the ordered pair (0,2) is a solution of -3x + y = -2 ....IT MEANS YOU SUBSTITUTE 0 FOR X AND 2 FOR Y IN THE GIVEN EQN.AND VERIFY WHETHER IT IS SATISFIED OR NOT...THAT IS IN THIS CASE FIND L.H.S.AND R.H.S. SEPERATELY AND CHECK WHETHER THEY COME OUT SAME OR NOT
L.H.S=-3X+Y=-3*0+2=0+2=2
R.H.S.=-2...
SO L.H.S IS NOT EQUAL TO R.H.S.
HENCE (0,2)IS NOT A SOLUTION OF THE GIVEN EQN.